# MNSC-0001-2024.00

### 15 Mar 2026 v1

@@ 7 Proofs of Theoretical Results in Section 4
According to the proof of Theorem 4.4, we observe that if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, then $x^{\gamma}_{i}=\tilde{x}_{i}^{\gamma}$ and $s(x^{\gamma}_{i})\coloneqq\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x^{\gamma}_{i},s)=\bar{s}_{i}$. Furthermore, since $(\partial/\partial x)g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is strictly decreasing on $0,\bar{x}_{i}]$ by Assumption 4.1(iv) and (21c), it suffices to show that $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $\bar{x}_{i}\leq\bar{s}_{i}$ imply
@@
[…]
Proof 7.5
Proof of Proposition 4.6. Fix $i\in[m]$ and $\bar{x}_{i}>0$. By Theorem 4.4, it suffices to prove that $x^{\gamma}_{i}$ decreases in $\gamma$ if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $\bar{s}_{i}$ is sufficiently small. Define $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s))$ for each $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$. Fix $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. We observe that if $\gamma>\sqrt{3}-1$, then for some $\theta\in[(\bar{x}_{i}+\bar{s}_{i})/2,(1+\gamma)(\bar{x}_{i}+\bar{s}_{i})/2]$,
@@
where $\delta=\min_{x\in[0,\bar{x}_{i}+\bar{s}_{i}]}Q^{\prime\prime}(x)>0$, the second equality follows from the first-order Taylor expansion, and the two inequalities hold since $Q^{\prime}(0)\leq Q^{\prime}((\bar{x}_{i}+\bar{s}_{i})/2)\leq Q^{\prime}((1+\gamma)(\bar{x}_{i}+\bar{s}_{i})/2)\leq 0$ by Assumption 4.1.
As in the proof of Theorem 4.4, define $\tilde{x}_{i}^{\gamma}$ as the solution of (11) with respect to $x$ given $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. In the remainder of the proof, we assume that $\bar{s}_{i}<-{(2-\sqrt{3})\bar{x}_{i}^{2}\delta}/({4Q^{\prime}(0)})$ and claim that
@@
This is trivial if $\tilde{x}_{i}^{\gamma}<(\bar{x}_{i}-\bar{s}_{i})/2$. Thus, suppose that $\tilde{x}_{i}^{\gamma}\geq(\bar{x}_{i}-\bar{s}_{i})/2$. Then, since $(\partial/\partial x)g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is decreasing on $0,\bar{x}_{i}]$ and $(\partial/\partial x)g_{i}^{\gamma}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})=0$ by the proof of Theorem 4.4, (33) and (57) suggests $\gamma\leq\sqrt{3}-1$ and
@@
[…]
8 Proofs of Theoretical Results in Section 5
@@
Proof 8.1
Proof of Lemma 5.1. To facilitate the proof, we use the following lemma, whose proof is provided in Appendix 9.
Lemma 8.2
@@
Since $\bm{\pi}(\cdot)$ is increasing and $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})=\bar{p}^{\bf A}$, Lemma 8.2 implies that
$\bar{\bm{\pi}}^{\bf A}=\bm{\pi}({\bf z}\bar{p}^{\bf A})\leq\bm{\pi}({\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A})=\bm{\pi}^{\bf A}.$
@@
$\psi_{j}^{\bf A}({\bf x})$ $=\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}_{j}^{\top}\bm{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$ (61)
$\leq\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}_{j}^{\top}\bar{\bm{\pi}}^{\bf A}\right)^{+}}{\bar{p}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$
$\leq\frac{\left(L_{j}-e_{j}-x_{j}\bar{p}^{\bf A}-{\bf A}_{j}^{\top}\bar{\bm{\pi}}^{\bf A}\right)^{+}}{\bar{p}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$
$=\left(\frac{L_{j}-e_{j}-{\bf A}_{j}^{\top}\bar{\bm{\pi}}^{\bf A}}{\bar{p}^{\bf A}}-x_{j}\right)\wedge\left(z_{j}-x_{j}\right)$
@@
[…]
Proof 8.3
Proof of Proposition 5.2. To explicitly show the dependence of $\bm{\pi}^{\bf A}$ on the first-period liability repayment $\bf x$}, we write $\bm{\pi}^{\bf A}=\bm{\pi}^{\bf A}_{\bf x}$ in this proof. We first show that
$\bm{\pi}^{\bf A}_{\bf x}\geq\bm{\pi}({\bf c}({\bf x}))~~\text{for all}~{\bf x}\leq\bar{{\bf x}}^{\bf A}.$ (62)
Define ${\bf\psi}^{\bf A}({\bf x})\coloneqq(\psi_{1}^{\bf A}({\bf x}),\dots,\psi_{m}^{\bf A}({\bf x}))^{\top}$. Let $\bm{\phi}_{0}={\bf L}\wedge\left({\bf e}+{\bf x}Q({\bf 1}^{\top}{\bf x})+(\bar{{\bf x}}^{\bf A}-{\bf x})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}^{\top}\bm{\pi}^{\bf A}_{\bf x}\right)$ and $\bm{\phi}_{k+1}={\bf L}\wedge\left({\bf e}+{\bf x}Q({\bf 1}^{\top}{\bf x})+(\bar{{\bf x}}^{\bf A}-{\bf x})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}^{\top}\bm{\phi}_{k}\right)\text{ for }k=0,1,2,\dots.$ Then, $\bm{\pi}^{\bf A}_{\bf x}\geq\bm{\phi}_{0}$ since for any $j\in[m]$,
$\pi_{{\bf x},j}^{\bf A}$ $=L_{j}\wedge\left(e_{j}+x_{j}Q({\bf 1}^{\top}{\bf x})+(z_{j}-x_{j})Q\big({\bf 1}^{\top}{\bf\psi}^{\bf A}({\bf x})\big)+{\bf A}_{j}^{\top}\bm{\pi}^{\bf A}_{\bf x}\right)$
$\geq L_{j}\wedge\left(e_{j}+x_{j}Q({\bf 1}^{\top}{\bf x})+(\bar{x}^{\bf A}_{j}-x_{j})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}_{j}^{\top}\bm{\pi}^{\bf A}_{\bf x}\right)$
@@
where the inequality follows from Lemma 5.1 and the fact that $\bar{{\bf x}}^{\bf A}\leq{\bf z}$. Therefore, $\bm{\pi}^{\bf A}_{\bf x}\geq\bm{\phi}_{0}$. Furthermore, using a similar argument, it can be shown that if $\bm{\pi}^{\bf A}_{\bf x}\geq\bm{\phi}_{k}$ for some $k\geq 0$, then $\bm{\pi}^{\bf A}_{\bf x}\geq\bm{\phi}_{k+1}$. Hence, $\bm{\pi}_{\bf x}^{\bf A}\geq\bm{\phi}_{k}$ for all $k$ by induction, which suggests (62) since $\bm{\phi}_{k}\rightarrow\bm{\pi}({\bf c}({\bf x}))$ as $k$ increases.
Next, we prove that $\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})\leq L_{i}$ for any $x_{i}\leq\bar{x}^{\bf A}_{i}$. Define a vector ${\bf u}\in\mathbb{R}^{m}$ with $u_{i}=x_{i}$ and ${\bf u}_{-i}=\bar{{\bf x}}^{\bf A}_{-i}$. Note that $\psi_{j}^{\bf A}({\bf u})=0$ for all $j\neq i$ by Lemma 5.1. If $\psi_{i}^{\bf A}({\bf u})>0$, then $L_{i}-e_{i}-x_{i}Q({\bf 1}^{\top}{\bf u})-{\bf A}^{\top}\bm{\pi}^{\bf A}_{\bf u}>0$ by (16), and hence,
$\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})$ $=e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+\psi_{i}^{\bf A}({\bf u})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\bm{\pi}^{\bf A}_{\bf u}$ (63)
$=\Big(L_{i}\vee\big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}^{\top}\bm{\pi}^{\bf A}_{\bf u}\big)\Big)\wedge\Big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+(z_{i}-x_{i})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\bm{\pi}^{\bf A}_{\bf u}\Big)$
$=L_{i}\wedge\Big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+(z_{i}-x_{i})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\bm{\pi}^{\bf A}_{\bf u}\Big)$
@@
where the second equality holds by (16). If $\psi_{i}^{\bf A}({\bf u})=0$, then $\bm{\pi}^{\bf A}_{\bf u}=\bm{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))$ by (15) and (16), and therefore,
$\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})$ $=e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}^{\top}\bm{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))$
$\leq e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}_{i}^{\top}\bm{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))-\left(e_{i}+\bar{x}^{\bf A}_{i}\bar{p}^{\bf A}+{\bf A}_{i}^{\top}\bar{\bm{\pi}}^{\bf A}\right)+L_{i}$
$=x_{i}Q({\bf 1}^{\top}{\bf u})-\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+{\bf A}_{i}^{\top}\big(\bm{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))-\bm{\pi}(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\big)+L_{i}$
$\leq x_{i}Q({\bf 1}^{\top}{\bf u})-\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+{\bf A}_{i}^{\top}\Big(\bm{\pi}\big(({\bf u}Q({\bf 1}^{\top}{\bf u}))\vee(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\big)-\bm{\pi}(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\Big)+L_{i}$
@@
where the first inequality comes from the definition of $\bar{x}^{\bf A}_{i}$ and the assumption that $\bar{{\bf x}}^{\bf A}>{\bf 0}$, the second inequality holds because $\bm{\pi}(\cdot)$ is increasing, the third inequality follows since $\bm{\pi}(\cdot)$ is non-expansive and ${\bf A}_{i}<{\bf 1}$, the third equality is satisfied as ${\bf u}_{-i}Q({\bf 1}^{\top}{\bf u})=\bar{{\bf x}}^{\bf A}_{-i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}^{\bf A}_{-i})\geq\bar{{\bf x}}^{\bf A}_{-i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})$ by Assumption 4.1(ii) and $u_{i}Q({\bf 1}^{\top}{\bf u})=x_{i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}^{\bf A}_{-i})\leq\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})$ by (21a), and the last inequality stems from Assumption 4.1(iii).
@@
$\psi_{i}^{\bf A}({\bf v})=\frac{\phi^{\bf A}_{i}(x_{i},{\bf v}_{-i}^{\bf A})-e_{i}-x_{i}Q\big({\bf 1}^{\top}{\bf v}\big)-{\bf A}_{i}^{\top}\bm{\pi}^{\bf A}_{\bf v}}{Q\left(\sum_{j=1}^{m}\psi_{j}^{\bf A}({\bf v})\right)}<\frac{\left(L_{i}-e_{i}-x_{i}Q\big({\bf 1}^{\top}{\bf v}\big)-{\bf A}_{i}^{\top}\bm{\pi}^{\bf A}_{\bf v}\right)^{+}}{Q\left(\sum_{j=1}^{m}\psi_{j}^{\bf A}({\bf v})\right)},$ (64)
@@
where the inequality holds by (62), Lemma 5.1, and the fact that $\bar{x}^{\bf A}_{i}\leq z_{i}$. Consequently, $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})<L_{i}$, which demonstrates the first statement by contrapositive.
[…]
Proof 8.4
Proof of Theorem 5.3. Fix $i\in[m]$. If $\bar{x}^{\bf A}_{i}=0$, the statement is straightforward. We assume that $\bar{x}^{\bf A}_{i}>0$. Note that $\Phi^{\bf A}_{i}(x,{\bf x}_{-i})=\Phi^{\bf A}_{i}(\bar{x}^{\bf A}_{i}-x,\bar{{\bf x}}^{\bf A}_{-i}-{\bf x}_{-i})$ for all $x\in[0,\bar{x}^{\bf A}_{i}]$. Thus, it suffices to show that for all $x\in[0,\bar{x}^{\bf A}_{i}/2)$,
@@
$\tilde{g}_{i}(y)\coloneqq\min_{\zeta\in{\cal Z}(y)}\min_{{\bf c}\in{\cal B}(y,\zeta)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\bm{\pi}({\bf c})\right\}=\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i}).$ (71)

### 17 Mar 2026 v2

@@ Appendix A Proofs of Theoretical Results in Section 4
According to the proof of Theorem 1, we observe that if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$, then $x^{\gamma}_{i}=\tilde{x}_{i}^{\gamma}$ and $s(x^{\gamma}_{i})\coloneqq\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x^{\gamma}_{i},s)=\bar{s}_{i}$. Furthermore, since $(\partial/\partial x)g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is strictly decreasing on $0,\bar{x}_{i}]$ by Assumption 1(iv) and (21c), it suffices to show that $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $\bar{x}_{i}\leq\bar{s}_{i}$ imply
@@
where the first and second inequalities hold by (55) and Assumption 1(v), respectively. ∎
Proof.
Proof of Proposition 2. Fix $i\in[m]$ and $\bar{x}_{i}>0$. By Theorem 1, it suffices to prove that $x^{\gamma}_{i}$ decreases in $\gamma$ if $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$ and $\bar{s}_{i}$ is sufficiently small. Define $g_{i}^{\gamma}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(1-\gamma)(x+s))$ for each $x\in[0,\bar{x}_{i}]$ and $s\in[0,\bar{s}_{i}]$. Fix $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. We observe that if $\gamma>\sqrt{3}-1$, then for some $\theta\in[(\bar{x}_{i}+\bar{s}_{i})/2,(1+\gamma)(\bar{x}_{i}+\bar{s}_{i})/2]$,
@@
where $\delta=\min_{x\in[0,\bar{x}_{i}+\bar{s}_{i}]}Q^{\prime\prime}(x)>0$, the second equality follows from the first-order Taylor expansion, and the two inequalities hold since $Q^{\prime}(0)\leq Q^{\prime}((\bar{x}_{i}+\bar{s}_{i})/2)\leq Q^{\prime}((1+\gamma)(\bar{x}_{i}+\bar{s}_{i})/2)\leq 0$ by Assumption 1.
As in the proof of Theorem 1, define $\tilde{x}_{i}^{\gamma}$ as the solution of (11) with respect to $x$ given $\gamma>\bar{s}_{i}/(\bar{x}_{i}+\bar{s}_{i})$. In the remainder of the proof, we assume that $\bar{s}_{i}<-{(2-\sqrt{3})\bar{x}_{i}^{2}\delta}/({4Q^{\prime}(0)})$ and claim that
@@
This is trivial if $\tilde{x}_{i}^{\gamma}<(\bar{x}_{i}-\bar{s}_{i})/2$. Thus, suppose that $\tilde{x}_{i}^{\gamma}\geq(\bar{x}_{i}-\bar{s}_{i})/2$. Then, since $(\partial/\partial x)g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is decreasing on $0,\bar{x}_{i}]$ and $(\partial/\partial x)g_{i}^{\gamma}(\tilde{x}_{i}^{\gamma},\bar{s}_{i})=0$ by the proof of Theorem 1, (33) and (57) suggests $\gamma\leq\sqrt{3}-1$ and
@@
where the first two inequalities comes from (59) and the third inequality holds by Assumption 1. Consequently, the above claim is demonstrated. This leads to the desired result because the proof of Theorem 1 shows that $(\partial/\partial x)g_{i}^{\gamma}(\cdot,\bar{s}_{i})$ is strictly decreasing on $0,\bar{x}_{i}]$ and that $x^{\gamma}_{i}=\tilde{x}_{i}^{\gamma}$ and $s(x^{\gamma}_{i})\coloneqq\mathop{\arg\min}_{s\in[0,\bar{s}_{i}]}g_{i}^{\gamma}(x^{\gamma}_{i},s)=\bar{s}_{i}$. ∎
Appendix B Proofs of Theoretical Results in Section 5
@@
Proof.
Proof of Lemma 2. To facilitate the proof, we use the following lemma, whose proof is provided in Appendix C.
Lemma 3.
@@
Since $\boldsymbol{\pi}(\cdot)$ is increasing and $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})=\bar{p}^{\bf A}$, Lemma 3 implies that
$\bar{\boldsymbol{\pi}}^{\bf A}=\boldsymbol{\pi}({\bf z}\bar{p}^{\bf A})\leq\boldsymbol{\pi}({\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A})=\boldsymbol{\pi}^{\bf A}.$
@@
$\psi_{j}^{\bf A}({\bf x})$ $=\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}_{j}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$ (61)
$\leq\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}_{j}^{\top}\bar{\boldsymbol{\pi}}^{\bf A}\right)^{+}}{\bar{p}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$
$\leq\frac{\left(L_{j}-e_{j}-x_{j}\bar{p}^{\bf A}-{\bf A}_{j}^{\top}\bar{\boldsymbol{\pi}}^{\bf A}\right)^{+}}{\bar{p}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)$
$=\left(\frac{L_{j}-e_{j}-{\bf A}_{j}^{\top}\bar{\boldsymbol{\pi}}^{\bf A}}{\bar{p}^{\bf A}}-x_{j}\right)\wedge\left(z_{j}-x_{j}\right)$
@@
where the second last equality holds since $x_{j}\leq\bar{x}^{\bf A}_{j}\leq(L_{j}-e_{j}-{\bf A}_{j}^{\top}\bar{\boldsymbol{\pi}}^{\bf A})/\bar{p}^{\bf A}$ for all $j$. This completes the proof. ∎
Proof.
Proof of Proposition 3. To explicitly show the dependence of $\boldsymbol{\pi}^{\bf A}$ on the first-period liability repayment $\bf x$}, we write $\boldsymbol{\pi}^{\bf A}=\boldsymbol{\pi}^{\bf A}_{\bf x}$ in this proof. We first show that
$\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\pi}({\bf c}({\bf x}))~~\text{for all}~{\bf x}\leq\bar{{\bf x}}^{\bf A}.$ (62)
Define ${\bf\psi}^{\bf A}({\bf x})\coloneqq(\psi_{1}^{\bf A}({\bf x}),\dots,\psi_{m}^{\bf A}({\bf x}))^{\top}$. Let $\boldsymbol{\phi}_{0}={\bf L}\wedge\left({\bf e}+{\bf x}Q({\bf 1}^{\top}{\bf x})+(\bar{{\bf x}}^{\bf A}-{\bf x})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf x}\right)$ and $\boldsymbol{\phi}_{k+1}={\bf L}\wedge\left({\bf e}+{\bf x}Q({\bf 1}^{\top}{\bf x})+(\bar{{\bf x}}^{\bf A}-{\bf x})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}^{\top}\boldsymbol{\phi}_{k}\right)\text{ for }k=0,1,2,\dots.$ Then, $\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\phi}_{0}$ since for any $j\in[m]$,
$\pi_{{\bf x},j}^{\bf A}$ $=L_{j}\wedge\left(e_{j}+x_{j}Q({\bf 1}^{\top}{\bf x})+(z_{j}-x_{j})Q\big({\bf 1}^{\top}{\bf\psi}^{\bf A}({\bf x})\big)+{\bf A}_{j}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf x}\right)$
$\geq L_{j}\wedge\left(e_{j}+x_{j}Q({\bf 1}^{\top}{\bf x})+(\bar{x}^{\bf A}_{j}-x_{j})Q\big({\bf 1}^{\top}(\bar{{\bf x}}^{\bf A}-{\bf x})\big)+{\bf A}_{j}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf x}\right)$
@@
where the inequality follows from Lemma 2 and the fact that $\bar{{\bf x}}^{\bf A}\leq{\bf z}$. Therefore, $\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\phi}_{0}$. Furthermore, using a similar argument, it can be shown that if $\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\phi}_{k}$ for some $k\geq 0$, then $\boldsymbol{\pi}^{\bf A}_{\bf x}\geq\boldsymbol{\phi}_{k+1}$. Hence, $\boldsymbol{\pi}_{\bf x}^{\bf A}\geq\boldsymbol{\phi}_{k}$ for all $k$ by induction, which suggests (62) since $\boldsymbol{\phi}_{k}\rightarrow\boldsymbol{\pi}({\bf c}({\bf x}))$ as $k$ increases.
Next, we prove that $\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})\leq L_{i}$ for any $x_{i}\leq\bar{x}^{\bf A}_{i}$. Define a vector ${\bf u}\in\mathbb{R}^{m}$ with $u_{i}=x_{i}$ and ${\bf u}_{-i}=\bar{{\bf x}}^{\bf A}_{-i}$. Note that $\psi_{j}^{\bf A}({\bf u})=0$ for all $j\neq i$ by Lemma 2. If $\psi_{i}^{\bf A}({\bf u})>0$, then $L_{i}-e_{i}-x_{i}Q({\bf 1}^{\top}{\bf u})-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}>0$ by (16), and hence,
$\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})$ $=e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+\psi_{i}^{\bf A}({\bf u})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}$ (63)
$=\Big(L_{i}\vee\big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}\big)\Big)\wedge\Big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+(z_{i}-x_{i})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}\Big)$
$=L_{i}\wedge\Big(e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+(z_{i}-x_{i})Q(\psi_{i}^{\bf A}({\bf u}))+{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf u}\Big)$
@@
where the second equality holds by (16). If $\psi_{i}^{\bf A}({\bf u})=0$, then $\boldsymbol{\pi}^{\bf A}_{\bf u}=\boldsymbol{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))$ by (15) and (16), and therefore,
$\phi^{\bf A}_{i}(x_{i},\bar{{\bf x}}^{\bf A}_{-i})$ $=e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}^{\top}\boldsymbol{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))$
$\leq e_{i}+x_{i}Q({\bf 1}^{\top}{\bf u})+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))-\left(e_{i}+\bar{x}^{\bf A}_{i}\bar{p}^{\bf A}+{\bf A}_{i}^{\top}\bar{\boldsymbol{\pi}}^{\bf A}\right)+L_{i}$
$=x_{i}Q({\bf 1}^{\top}{\bf u})-\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+{\bf A}_{i}^{\top}\big(\boldsymbol{\pi}({\bf u}Q({\bf 1}^{\top}{\bf u}))-\boldsymbol{\pi}(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\big)+L_{i}$
$\leq x_{i}Q({\bf 1}^{\top}{\bf u})-\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})+{\bf A}_{i}^{\top}\Big(\boldsymbol{\pi}\big(({\bf u}Q({\bf 1}^{\top}{\bf u}))\vee(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\big)-\boldsymbol{\pi}(\bar{{\bf x}}^{\bf A}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A}))\Big)+L_{i}$
@@
where the first inequality comes from the definition of $\bar{x}^{\bf A}_{i}$ and the assumption that $\bar{{\bf x}}^{\bf A}>{\bf 0}$, the second inequality holds because $\boldsymbol{\pi}(\cdot)$ is increasing, the third inequality follows since $\boldsymbol{\pi}(\cdot)$ is non-expansive and ${\bf A}_{i}<{\bf 1}$, the third equality is satisfied as ${\bf u}_{-i}Q({\bf 1}^{\top}{\bf u})=\bar{{\bf x}}^{\bf A}_{-i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}^{\bf A}_{-i})\geq\bar{{\bf x}}^{\bf A}_{-i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})$ by Assumption 1(ii) and $u_{i}Q({\bf 1}^{\top}{\bf u})=x_{i}Q(x_{i}+{\bf 1}^{\top}\bar{{\bf x}}^{\bf A}_{-i})\leq\bar{x}^{\bf A}_{i}Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})$ by (21a), and the last inequality stems from Assumption 1(iii).
@@
$\psi_{i}^{\bf A}({\bf v})=\frac{\phi^{\bf A}_{i}(x_{i},{\bf v}_{-i}^{\bf A})-e_{i}-x_{i}Q\big({\bf 1}^{\top}{\bf v}\big)-{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf v}}{Q\left(\sum_{j=1}^{m}\psi_{j}^{\bf A}({\bf v})\right)}<\frac{\left(L_{i}-e_{i}-x_{i}Q\big({\bf 1}^{\top}{\bf v}\big)-{\bf A}_{i}^{\top}\boldsymbol{\pi}^{\bf A}_{\bf v}\right)^{+}}{Q\left(\sum_{j=1}^{m}\psi_{j}^{\bf A}({\bf v})\right)},$ (64)
@@
where the inequality holds by (62), Lemma 2, and the fact that $\bar{x}^{\bf A}_{i}\leq z_{i}$. Consequently, $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}_{-i}]}\Phi^{\bf A}_{i}(x_{i},{\bf x}_{-i})<L_{i}$, which demonstrates the first statement by contrapositive.
The second statement is straightforward if $\phi^{\bf A}_{i}(x_{i},{\bf v}_{-i}^{\bf A})=L_{i}$. Otherwise, it holds by (65). ∎
Proof.
Proof of Theorem 3. Fix $i\in[m]$. If $\bar{x}^{\bf A}_{i}=0$, the statement is straightforward. We assume that $\bar{x}^{\bf A}_{i}>0$. Note that $\Phi^{\bf A}_{i}(x,{\bf x}_{-i})=\Phi^{\bf A}_{i}(\bar{x}^{\bf A}_{i}-x,\bar{{\bf x}}^{\bf A}_{-i}-{\bf x}_{-i})$ for all $x\in[0,\bar{x}^{\bf A}_{i}]$. Thus, it suffices to show that for all $x\in[0,\bar{x}^{\bf A}_{i}/2)$,
@@
$\tilde{g}_{i}(y)\coloneqq\min_{\zeta\in{\cal Z}(y)}\min_{{\bf c}\in{\cal B}(y,\zeta)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c})\right\}=\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i}).$ (71)